HIGH RESOLUTION FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE Ã^{2}Π − X̃^{2}Π TRAN-SITION OF THE OCS^{+} ION

\underline{YOSHIHIRO NAKASHIMA}, KEIICHI TANAKA, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan..

The OCS^{+} ion was produced by Penning ionization of OCS with metastable He^{∗}, and UV emission in the 370 − 500 nm region was measured by Fourier transform spectrometer. The Ã^{2}Π − X̃^{2}Π electronic transition has been reported in the 300 − 500 nm region, but only the emission bands from the vibrational ground state of the Ã^{2}Π state were observed because of the predissociation in the Ã^{2}Π state ^{a}. High resolution study of the 0-0 band has been reported for both the Ω = 3/2 and Ω = 1/2 spin states ^{b}^{,} ^{c}. In the present study, seven rotationally resolved bands were observed in the 370 − 500 nm region and assigned to the CO stretch (ν′′_{3}) progression : Ã^{2}Π_{3/2}(000) − X̃^{2}Π_{3/2}(00v) (v = 2 − 5) and Ã^{2}Π_{1/2}(000) − X̃^{2}Π_{1/2}(00v) (v = 3 − 4). The observed lines of the Ω = 1/2 substate were split into the two components due to the Λ-type doubling. The spectra for Ω = 3/2 and Ω = 1/2 were analysed separately to obtain molecular constants for each spin state. Band origins, vibrational frequencies, rotational constants, vibrational-rotational constants for both spin substates and Λ-type doubling constant for the Ω = 1/2 substate were determined by a least squares fitting. Rotational constants in the vibrational ground state were determined for both spin states. The spin-orbit interaction constants derived from these rotational constants, A′ = −125.0 cm^{−1} and A′′ = −383.8 cm^{−1}, are consistent with those derived from the photodissociation study ^{d}.

aD. L. Judge and L. C. Lee, Int. J. Mass Spectrom. Ion Phys., 17, 329, (1975)

bM. Oschner, M. Tsuji, and J. P. Maier, Chem. Phys. Lett., 115, 373, (1985)

cR. Weinkauf and U. Boesl, J. Chem. Phys., 101, 8482, (1994)

dR. Kakoschke, U. Boesl, J. Hermann and E. W. Schlag, Chem. Phys. Lett., 119, 467, (1985)